OpenStudy (anonymous):

f(x)=2x^2-x-1 if f(x)=-1, what is x? What point(s) are on the graph of f?

OpenStudy (anonymous):

its easy . f(x)=2x^2-x-1 now f(x)=-1 by plugging values we get -1=2x^2-x-1 which is equal to 0=2x^2-x take x as common we get 0=x(2x-1) so x=0 or x=1/2 now we can find points